Logiweb dictionary of prop

Up Help

0 0 prop

1 2 ∗ ⇒ ∗ 2 0 A1 3 0 A2 4 0 A3 5 0 MP 6 0 PropositionalCalculus 7 0 AutoImply

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-06-02.UTC:08:43:13.701947 = MJD-53523.TAI:08:43:45.701947 = LGT-4624418625701947e-6